OpenStudy (anonymous):

-2x-1=(3x+5)/(2) The answer is -1 Can somebody show me the steps

OpenStudy (anonymous):

\[2(-2x-1)=3x+5\] \[-4x-2=3x+5\] \[-7x=7\] \[x=-1\]